# Modeling AdaGrad, RMSProp, and Adam with Integro-Differential Equations

### v1

@@ Abstract
In this paper, we propose a continuous-time formulation for the AdaGrad, RMSProp, and Adam optimization algorithms by modeling them as first-order integro-differential equations. We perform numerical simulations of these equations to demonstrate their validity as accurate approximations of the original algorithms. Our results indicate a strong agreement between the behavior of the continuous-time models and the discrete implementations, thus providing a new perspective on the theoretical understanding of adaptive optimization methods.
@@ 1 Introduction
While Stochastic Gradient Descent (SGD) Bottou (2010) extends the Gradient Descent method by using mini-batches, or randomly selected subsets of data, to compute gradients, this article focuses on a non-stochastic perspective. The continuous nature of these methods permits a more direct application of differential equation techniques. For readers interested in a continuous description of the stochastic method, we refer to Sirignano & Spiliopoulos (2017).
@@
  • We provide a numerical comparison between the results obtained from these continuous models and the original discrete algorithms, focusing on the accuracy and dynamics of the continuous approximations: Figures 8 to 14. Additionally, we detail the numerical methods employed to solve the integro-differential equations.
@@
    Optimization Algorithms. We review the AdaGrad, RMSProp, and Adam algorithms and present numerical simulations that illustrate their dynamic behavior.
@@ 1.1 Notation
In this subsection, we introduce and define the notation used throughout the article. Although the notation may appear “highly” physical, it is selected for its simplicity and to enhance the reader’s comprehension.
@@
    $\nabla_{\theta}=\left(\frac{\partial}{\partial\theta^{1}},\ldots,\frac{\partial}{\partial\theta^{i}},\ldots,\frac{\partial}{\partial\theta^{n}}\right):=\left(\partial_{1},\ldots,\partial_{i},\ldots,\partial_{n}\right).$
@@ 1.2 Related Work
2 Optimization Algorithm
In this section, we present the optimization algorithms [^1] For a comprehensive overview of Gradient Descent optimization algorithms, see Ruder (2016).: AdaGrad, RMSProp, and Adam, with the aim of illustrating the evolution of the parameters as they converge towards the minimum of the function:
$f(\theta)=(\theta-4)^{2}.$
This function is chosen for its convexity and differentiability, with a known minimum at $\theta=4$. The simplicity of this function allows for straightforward analytical determination of the minimum, facilitating direct comparisons with numerical simulations. For an additional example involving the Mean Squared Error (MSE) loss function, refer to the appendix B.
@@ 2.1 Fundamentals of AdaGrad
AdaGrad is distinguished by maintaining a cumulative sum of the squares of past gradients for each parameter, denoted as $G_{k}$. This accumulation effectively scales the learning rate inversely proportional to the square root of $G_{k}$, ensuring that parameters frequently updated receive smaller updates, while parameters updated less frequently receive larger updates.
@@
2.1.1 Mathematical Formulation of AdaGrad and Numerical Simulations
@@
Numerical Simulation:
Figure 1 illustrates the convergence trajectories of $\theta$ as the AdaGrad algorithm minimizes the function $f(\theta)=(\theta-4)^{2}$, using two different learning rates ($\alpha=0.1$ and $\alpha=0.01$). For $\alpha=0.1$, $\theta$ converges towards the minimum value of 4 within fewer than 2000 iterations. The curve flattens as the algorithm stabilizes at the minimum. Conversely, with a smaller learning rate of $\alpha=0.01$, convergence is significantly slower, taking over 150,000 iterations to reach the same minimum. This is due to the convex nature of the function, where larger steps allow the algorithm to reach the minimum more quickly compared to smaller steps.
Figure 1: Trajectory of the Discrete AdaGrad Algorithm. The plot shows the convergence of $\theta$ towards the minimum of the function for two learning rates ($\alpha=0.1$ and $\alpha=0.01$). In both cases, $\theta$ converges to 4, the minimum of the function $f(\theta)=(\theta-4)^{2}$.
Figure 2 shows the accumulation of squared gradients $G_{k}$ over the iterations for the same learning rates. With $\alpha=0.1$, the accumulation of gradients stabilizes relatively quickly, corresponding to the fast convergence observed in Figure 1. For $\alpha=0.01$, however, the accumulation grows more gradually, reflecting the slower convergence of $\theta$.
Figure 2: Trajectory of the Accumulated Gradient for the AdaGrad Algorithm. The plot tracks the value of $G_{k}$ as it is iteratively updated by the AdaGrad optimizer to converge towards the minimum of the function $f(\theta)=(\theta-4)^{2}$.
@@ 2.2 Fundamentals of RMSProp
2.2.1 Mathematical Formulation of RMSProp and Numerical Simulations
@@
Numerical Simulations:
Figure 3 illustrates the convergence trajectories of $\theta$ for different values of $\beta\in\{0.0,0.9,0.99\}$ and learning rates ($\alpha=0.1$ and $\alpha=0.01$). With a higher learning rate of $\alpha=0.1$, RMSProp exhibits varied behavior depending on the value of $\beta$:
For $\beta=0.0$, the algorithm converges quickly but shows significant oscillations near the minimum, indicating a high sensitivity to recent gradients and leading to unstable updates. As $\beta$ increases to 0.9 and 0.99, the oscillations diminish, resulting in smoother convergence. Furthermore, with $\beta=0.99$, the convergence speed improves, demonstrating a positive correlation between stability and convergence rate.
When the learning rate is reduced to $\alpha=0.01$, the overall convergence is slower but remains stable across all values of $\beta$. Higher values of $\beta$ continue to offer a smoother trajectory toward the minimum.
Figure 3: Trajectory of the Discrete RMSProp Algorithm. The plot illustrates the convergence trajectories of $\theta$ for different values of $\beta\in\{0.0,0.9,0.99\}$ across two learning rates ($\alpha=0.1$ and $\alpha=0.01$). In all cases, $\theta$ converges towards the minimum value of 4, corresponding to the minimum of the function $f(\theta)=(\theta-4)^{2}$. Compared to AdaGrad, RMSProp converges faster, but oscillations can occur when $\beta$ is set to or close to 0. These oscillations diminish as $\beta$ increases, resulting in smoother convergence.
Figure 4 compares the behavior of the squared gradients $v_{k}$ for different values of $\beta\in\{0.0,0.9,0.99\}$. In RMSProp, the squared gradients tend to saturate at 0.0, a behavior that contrasts with AdaGrad, where the squared gradients stabilize at non-zero values.
With $\alpha=0.1$, RMSProp demonstrates that the squared gradients $v_{k}$ decrease rapidly when $\beta=0.0$, often leading to saturation at 0.0. As $\beta$ increases, the decrease in $v_{k}$ becomes more gradual, reducing the likelihood of immediate saturation. At a lower learning rate of $\alpha=0.01$, the squared gradients in RMSProp decrease more slowly, particularly with higher values of $\beta$, where the initial bump becomes more pronounced but ultimately still declines to 0.
Figure 4: Trajectory of the Accumulated Gradient for the RMSProp Algorithm. The plot tracks the convergence of the accumulated squared gradients $v_{k}$ for the same $\beta$ and $\alpha$ configurations shown in Figure 3. In all cases, the squared gradients $v_{k}$ decrease over time as the optimizer progresses. Unlike AdaGrad, where squared gradients stabilize at non-zero values, RMSProp tends to saturate at 0.0 as the learning rate decreases. Depending on the value of $\beta$, the squared gradients either decay directly to 0 or exhibit an initial bump that quickly falls to 0.
@@ 2.3 Fundamentals of Adam
2.3.1 Mathematical Formulation of Adam and Numerical Simulations
@@
1:Initialize the parameters $\theta^{i}_{k=0}$, the first moment $m^{i}_{k=0}=0$, and the second moment $v^{i}_{k=0}=0$.
@@
Numerical Simulations:
Figure 5 illustrates the convergence of $\theta$ to the minimum, $\theta=4$, of the objective function $f(\theta)=(\theta-4)^{2}$ using the Adam optimizer, under different configurations [^2] The motivation for choosing this configuration of $\beta$ is given by the article Kingma & Ba (2014). of $\beta_{1}\in\{0.0,0.9\}$ and $\beta_{2}\in\{0.99,0.999\}$. With a learning rate of $\alpha=0.1$, $\theta$ converges rapidly to the minimum, particularly when both $\beta_{1}$ and $\beta_{2}$ are high (0.9 and 0.99). However, small oscillations around the minimum are observed when $\beta_{1}=0.9$, indicating slight instability in convergence. As $\beta_{1}$ decreases, these oscillations diminish, allowing for smoother and more stable convergence to the minimum value. At a lower learning rate of $\alpha=0.01$, $\theta$ converges more slowly but steadily towards the minimum. Configurations with high $\beta_{1}$ and $\beta_{2}$ show more controlled convergence, with no oscillations. Compared to AdaGrad and RMSProp, Adam offers a balanced approach: it converges faster than AdaGrad but is slightly slower than RMSProp. Both Adam and AdaGrad or RMSPRop share the characteristic that, as the learning rate decreases, convergence to the minimum takes longer, as discussed earlier.
Figure 5: Trajectory of the Discrete Adam Algorithm. The plot shows how $\theta$ converges to the minimum ($\theta=4$) of the objective function $f(\theta)=(\theta-4)^{2}$ using the Adam optimizer under different configurations of $\beta_{1}$ and $\beta_{2}$. Higher values of $\beta_{1}$ and $\beta_{2}$ generally result in fewer oscillations around the minimum. These oscillations tend to disappear as the learning rate decreases.
Figure 6 illustrates the convergence trajectories of the first moment $m^{i}_{k}$ for the Adam optimizer, considering different values of $\beta_{1}$ and $\beta_{2}$, under two learning rates ($\alpha=0.1$ and $\alpha=0.01$). With a learning rate of $\alpha=0.1$, it is evident that $m_{k}$ exhibits significant initial oscillations, particularly when $\beta_{1}=0.9$. As $\beta_{1}$ decreases to 0.0, these oscillations diminish, resulting in smoother updates to the moving averages. Regardless of the values of $\beta_{1}$, the moment $m^{i}_{k}$ stabilizes around the same iteration. In contrast, at a lower learning rate of $\alpha=0.01$, the oscillations in $m_{k}$ disappear, leading to a more stable convergence. In all cases, $m_{k}$ stabilizes around 0.
Figure 7 presents the convergence trajectories of the second moment $v_{k}$ under the Adam optimizer, across different configurations of $\beta_{1}$, $\beta_{2}$, and $\alpha$. With a learning rate of $\alpha=0.1$, $v_{k}$ shows significant initial variation, especially when $\beta_{2}=0.99$. Here, $v_{k}$ increases rapidly and then decreases as the algorithm progresses. When $\alpha=0.01$, a similar dynamic is observed, but the peak is more pronounced for $\beta_{2}=0.99$, while for $\beta_{2}=0.999$, it rises slightly higher. The initial behavior of $v_{k}$ in Adam resembles that of RMSProp, with an initial peak followed by a smooth decline. However, not all configurations stabilize at 0.0 when the minimum is reached.
Figure 6: Trajectory of the First Moment for Adam Algorithm. The convergence trajectories of the first moment $m_{k}$ for different values of $\beta_{1}$ and $\beta_{2}$ at two learning rates ($\alpha=0.1$ and $\alpha=0.01$). Higher learning rates ($\alpha=0.1$) show more pronounced oscillations, especially when $\beta_{1}$ is high, which diminish as $\beta_{1}$ decreases. Lower learning rates ($\alpha=0.01$) result in more stable trajectories.
Figure 7: Trajectory of the Second Moment for Adam Algorithm. The convergence trajectories of the second moment $v_{k}$ under different configurations of $\beta_{1}$, $\beta_{2}$, and $\alpha$. Higher values of $\beta_{1}$ and $\beta_{2}$ result in a more gradual increase and smoother stabilization of $v_{k}$, while lower values lead to sharper peaks and rapid decreases.
3 Time-Continuous models for AdaGrad, RMSProp and Adam
@@
Under Assumption (2) and with an initial value for the accumulated gradients $G_{0}=0$, the continuous nonlocal dynamics of AdaGrad can be characterized by the following equation:
@@
Appendix A. ∎
@@
Under Assumption (2) and given an initial value for the squared gradient moving average $v_{0}=0$, the continuous nonlocal dynamics of RMSProp can be characterized by the following equation:
@@
Appendix A. ∎
@@
This modification aligns with the behavior observed in Figure 4, where an exponential decay is evident as time (or iterations, in this context) progresses.
@@
Under Assumption (2) and with initial values for the moments $m^{i}_{0}=0$ and $v_{0}=0$, the Adam optimizer can be characterized by the following continuous nonlocal dynamics:
@@
Appendix A. ∎
In the case of Adam, the nonlocality is encapsulated within both moments, as expected. A key distinction between Adam and other optimizers like AdaGrad or RMSProp (beyond the use of the exponential decay rates, $\beta_{1}$ and $\beta_{2}$, and the specific kernel structure in AdaGrad) is that Adam does not require incorporating the future time differential $\alpha$ that was necessary in the other cases.
@@
Extensions to modifications of these algorithms, such as AdamW or AdamL2, can be derived similarly by adding:
  • For AdamW:

### v2

@@ Abstract
In this paper, we propose a continuous-time formulation for the AdaGrad, RMSProp, and Adam optimization algorithms by modeling them as first-order integro-differential equations. We perform numerical simulations of these equations, along with stability and convergence analyses, to demonstrate their validity as accurate approximations of the original algorithms. Our results indicate a strong agreement between the behavior of the continuous-time models and the discrete implementations, thus providing a new perspective on the theoretical understanding of adaptive optimization methods.
Keywords: Adaptive Optimization Algorithms, Integro-differential Equations, Machine Learning.
@@ 1 Introduction
While Stochastic Gradient Descent (SGD) Bottou (2010) extends the Gradient Descent method by using mini-batches, or randomly selected subsets of data, to compute gradients, this article focuses on a non-stochastic perspective. The continuous nature of these methods permits a more direct application of differential equation techniques, as we demonstrate in Section 4. For readers interested in a continuous description of the stochastic method, we refer to Sirignano & Spiliopoulos (2017).
@@
  • We prove the stability of these integro-differential equations, as well as their convergence rates. Theorems 1 to 6.
  • We provide a numerical comparison between the results obtained from these continuous models and the original discrete algorithms, focusing on the accuracy and dynamics of the continuous approximations: Figures 1 to 15. Additionally, we detail the numerical methods employed to solve the integro-differential equations.
@@
    Optimization Algorithms. We review the AdaGrad, RMSProp, and Adam algorithms.
@@
    Convergence and Stability Analysis. We establish stability and convergence theorems for these continuous nonlocal models
  4. Sec. 5.
@@ 1.1 Notation
In this subsection, we introduce and define the notation used throughout Section  2 and 3. Although the notation may appear “highly” physical, it is adopted here for its simplicity and ease when presenting the results.
@@
    $\nabla_{\theta}=\left(\nabla_{1},\ldots,\nabla_{i},\ldots,\nabla_{n}\right)=\left(\frac{\partial}{\partial\theta^{1}},\ldots,\frac{\partial}{\partial\theta^{i}},\ldots,\frac{\partial}{\partial\theta^{n}}\right):=\left(\partial_{1},\ldots,\partial_{i},\ldots,\partial_{n}\right).$
In Section 4, we use the traditional notation used in optimization textbooks, where $\langle\cdot,\cdot\rangle$ denotes the scalar product in the $n$-dimensional Euclidean space, as we wish to adhere as closely as possible to standard mathematical notation.
@@ 1.2 Related Work
2 Optimization Algorithms
In this section, we present the optimization algorithms [^1] For a comprehensive overview of Gradient Descent optimization algorithms, see Ruder (2016).: AdaGrad, RMSProp, and Adam, with the aim of illustrating the evolution of the parameters as they converge towards the minimum of two representative toy models. The first model is a convex and differentiable function,
$f(\theta)=(\theta-4)^{2},$
chosen for its simplicity and for having a known minimum at $\theta=4$, which allows for straightforward analytical determination and facilitates direct comparisons with numerical simulations.
To complement this, we also consider a non-convex toy model,
$f(\theta)=\frac{1}{4}(\theta^{2}-1)^{2},$
which exhibits multiple local minima, $\theta=\pm 1$, and an unstable critical point (a local maxima) at $\theta=0$. With these two toy models, we can analyze the dynamics of the continuous-time formulations in both convex and non-convex regimes.
@@ 2.1 Fundamentals of AdaGrad
AdaGrad is distinguished by maintaining a cumulative sum of the squares of past gradients for each parameter, denoted as $G_{k,i}$. This accumulation effectively scales the learning rate inversely proportional to the square root of $\text{diag}(G_{k,i})$, ensuring that parameters frequently updated receive smaller updates, while parameters updated less frequently receive larger updates. This particular variant of AdaGrad is known as AdaGrad-Diagonal. However, for simplicity of analysis, in this work we focus on the AdaGrad-Norm variant, in which all parameters are scaled uniformly, namely, $G_{k}$.
@@
Mathematical Formulation of AdaGrad:
@@ 2.2 Fundamentals of RMSProp
Mathematical Formulation of RMSProp:
@@ 2.3 Fundamentals of Adam
Mathematical Formulation of Adam:
@@
1:Initialize the parameters $\theta^{i}_{k=0}$, the first moment $m^{i}_{k=0}=0$, and the second moment $v_{k=0}=0$.
@@
3 Continuous-time models for AdaGrad, RMSProp and Adam
@@
Under Assumptions 1 and 2, and with an initial value for the accumulated gradients $G_{0}=0$, the continuous nonlocal dynamics of AdaGrad can be characterized by the following equation:
@@
By applying Assumption 2, the evolution of the accumulated gradients can be described by the differential equation:
whose solution (for $t>0$) is:
Using Assumption 2 again, we derive the integro-differential equation that governs the updated parameters:
@@
Under Assumptions 1 and 2, and given an initial value for the squared gradient moving average $v_{0}=0$, the continuous nonlocal dynamics of RMSProp can be characterized by the following equation:
@@
By applying Assumption 2, the evolution of the squared gradient moving average is governed by the differential equation:
$\dot{v}+\frac{1-\beta}{\alpha}v=\frac{1-\beta}{\alpha}\partial^{i}f(\theta)\partial_{i}f(\theta)\,.$ (5)
Rewriting this equation yields:
$\frac{{\rm d}}{{\rm d}t}\left[e^{\frac{1-\beta}{\alpha}t}v\right]=\frac{1-\beta}{\alpha}e^{\frac{1-\beta}{\alpha}t}\partial^{i}f(\theta)\partial_{i}f(\theta)\,$, (6)
which leads to the solution (for $t>0$):
Finally, by applying Assumption 2, the parameter updates follow the integro-differential equation:
@@
Under Assumption 2 and 1, and with initial values for the moments $m^{i}_{0}=0$ and $v_{0}=0$, the Adam optimizer can be characterized by the following continuous nonlocal dynamics:
@@
Applying Assumptions 2 and 1 and performing some algebraic manipulations yields:
where
In the case of Adam, the nonlocality is encapsulated within both moments, as expected. A key distinction between Adam and other optimizers such as AdaGrad or RMSProp (beyond the use of the exponential decay rates, $\beta_{1}$ and $\beta_{2}$, and the specific kernel structure in AdaGrad) is that Adam does not require incorporating the future time differential $\alpha$ that was necessary in the other cases; namely, it is evaluated at the same instant $t$, so the nonlocality enters only through $m$ and $v$, without any shift to $t+\alpha$.
@@
4 Theoretical Framework for Continuous-Time Dynamics
Thanks to the continuous-time formulation of these algorithms, we can investigate their convergence and stability properties using mathematical tools such as Lyapunov or LaSalle stability analysis and convergence theory for both convex and non-convex objective functions. To enhance clarity, this section is divided into two parts: the first focuses on convex functions, while the second addresses non-convex functions.
4.1 General Assumptions
To begin, we assume that $f:\mathbb{R}^{n}\to\mathbb{R}$ is of class $\mathcal{C}^{1}$. We denote by $\theta(t)\in\mathbb{R}^{n}$ the continuous-time trajectory of the parameter, and by $g_{i}(t):=\nabla_{i}f(\theta(t))$ the gradient evaluated along this trajectory. Furthermore, in order to represent compactly the kernels of the integral operators appearing in the integro-differential equations, we introduce the following definition.
Definition 1.
Let $\nu$ be a finite Borel measure on $0,\infty)$. We define the kernel function
$K_{\nu}(t):=\int_{[0,\infty)}e^{-\lambda t}\,\mathrm{d}\nu(\lambda),\qquad t\geq 0.$ (15)
Given a function $g:[0,\infty)\to\mathbb{R}^{n}$ that is measurable and locally integrable, the associated memory operator is defined as
$M_{\nu,i}[g](t):=\int_{0}^{t}K_{\nu}(t-\tau)\,g_{i}(\tau)\,\mathrm{d}\tau.$ (16)
Lemma 1.
Let $\alpha>0$ and $\beta_{a}\in[0,1)$. If
$\nu_{\mathrm{AG}}=\frac{1}{\alpha}\,\delta_{0},\qquad\nu_{\mathrm{RM/Adam}}=\frac{1-\beta_{a}}{\alpha}\,\delta_{\frac{1-\beta_{a}}{\alpha}},$ (17)
then the corresponding kernels defined in Definition 1 are
$K_{\nu_{\mathrm{AG}}}(t)=\frac{1}{\alpha},\qquad K_{\nu_{\mathrm{RM/Adam}}}(t)=\frac{1-\beta_{a}}{\alpha}e^{-\frac{1-\beta_{a}}{\alpha}t},$ (18)
where $\delta_{a}$ denotes the Dirac measure, i.e.,
$\delta_{a}(B)=\begin{cases}1,&\text{if }a\in B,\\ 0,&\text{otherwise}.\end{cases}$ (19)
Proof.
Straightforward calculation. ∎
Next, we introduce the necessary lemmas, propositions, and regularity assumptions that will serve as the basis for proving more specific results —both lemmas and theorems—for convex and non-convex functions.
Lemma 2 (Bounds on $K_{\nu}$).
For all $t\geq 0$,
$0\ \leq\ K_{\nu}(t)\ \leq\ \|\nu\|,\quad\text{and in particular}\quad K_{\nu}(0)=\|\nu\|.$
Proof.
Since $e^{-\lambda t}\in[0,1]$ for all $\lambda\geq 0$ and $t\geq 0$, we have
[…]
The identity $K_{\nu}(0)=\|\nu\|$ follows immediately by taking $t=0$ in the definition of $K_{\nu}$. For AdaGrad, this yields $\|\nu_{\mathrm{AG}}\|=1/\alpha$, and for RMSProp/Adam, $\|\nu_{\mathrm{RM/Adam}}\|=(1-\beta_{a})/\alpha$. ∎
Proposition 4 (Bounds on the Integral of $K_{\nu}$).
For all $h\geq 0$,
$\int_{0}^{h}K_{\nu}(\sigma)\,\mathrm{d}\sigma\ \leq\ \|\nu\|\,h.$
Proof.
From Lemma 2 we have $K_{\nu}(\sigma)\leq\|\nu\|$ for all $\sigma\geq 0$. Integrating over $[0,h]$ yields
$\int_{0}^{h}K_{\nu}(\sigma)\,\mathrm{d}\sigma\ \leq\ \int_{0}^{h}\|\nu\|\,\mathrm{d}\sigma\ =\ \|\nu\|\,h.$
Lemma 3.
If $\nu$ is a finite Borel measure on $0,\infty)$, then $K_{\nu}$ is completely monotone:
$\frac{\mathrm{d}^{k}}{\mathrm{d}t^{k}}K_{\nu}(t)=(-1)^{k}\int\lambda^{k}e^{-\lambda t}\,\mathrm{d}\nu(\lambda),\qquad\forall k\in\mathbb{N}\,.$
In particular, it is non-increasing since $\dot{K}_{\nu}(t)\leq 0$ for all $t\geq 0$.
Proof.
This follows directly from differentiation under the integral sign and the positivity of $\nu$. ∎
Assumption 3 ($L$-smoothness).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be of class $C^{1}$. There exists $L>0$ such that
$\|\nabla f(x)-\nabla f(y)\|\ \leq\ L\,\|x-y\|\qquad\forall\,x,y\in\mathbb{R}^{n},$
equivalently,
$f(y)\ \leq\ f(x)+\langle\nabla f(x),\,y-x\rangle+\tfrac{L}{2}\|y-x\|^{2}\qquad\forall\,x,y\in\mathbb{R}^{n}.$
To conclude the general lemmas and propositions,
Definition 2.
Throughout, the notation $M_{\nu}[g^{2}](t)$ is understood as
$M_{\nu}[\|g\|^{2}](t)\ :=\ \int_{0}^{t}K_{\nu}(t-\tau)\,\|g(\tau)\|^{2}\,\mathrm{d}\tau.$
This convention avoids ambiguities that could arise from a component-wise interpretation.
Lemma 4 (Lower bound induced by $\epsilon$).
For all $t\geq 0$,
$\sqrt{M_{\nu}[g^{2}](t)+\epsilon}\;\geq\;\sqrt{\epsilon}.$ (20)
Proof.
Since the kernel $K_{\nu}(t)\geq 0$, it follows that $M_{\nu}[g^{2}](t)\geq 0$. Adding $\epsilon>0$ to both sides yields the claimed inequality immediately. ∎
With these definitions in place, we are equipped to address the convex and non-convex cases. Nevertheless, Table (1) provides a useful clarification for the reader, as it highlights the correspondence between the notation introduced in this section and that employed in the preceding one. The primary motivation for adopting this notation (in this section) is its suitability for the generalization of the theorems.
Section 3 Section 4
$G(t)$ $M_{\nu_{\text{AG}}}[g^{2}](t)$
$v(t)$ $M_{\nu_{\text{RM/Adam}}}[g^{2}](t)$
$m^{i}(t)$ $M^{i}_{\nu_{\text{Adam}}}[g](t)$
Table 1: Correspondence between the notation used in Section 3 and Section 4.
4.2 Case I: Convex Objective Functions
In this subsection, we shall concentrate on convex objective functions, under the assumptions of regularity —Assumption 3— and convexity Bauschke & Combettes (2017).
Assumption 4 (Strong convexity).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be of class $C^{1}$. There exists $m>0$ such that
$\langle\nabla f(x)-\nabla f(y),\,x-y\rangle\ \geq\ m\,\|x-y\|^{2}\qquad\forall\,x,y\in\mathbb{R}^{n}.$
equivalently,
$f(y)\ \geq\ f(x)+\langle\nabla f(x),\,y-x\rangle+\tfrac{m}{2}\|y-x\|^{2}\qquad\forall\,x,y\in\mathbb{R}^{n},$
Lemma 5.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be $m$-strongly convex and $L$-smooth, with minimizer $x^{*}$. Then, for all $x\in\mathbb{R}^{n}$,
$\frac{1}{2L}\,\|\nabla f(x)\|^{2}\ \leq\ f(x)-f(x^{*})\ \leq\ \frac{1}{2m}\,\|\nabla f(x)\|^{2}.$
Proof.
(a) Lower bound. If $f$ is $L$-smooth, then for all $x,y\in\mathbb{R}^{n}$,
$f(y)\ \leq\ f(x)+\langle\nabla f(x),\,y-x\rangle+\frac{L}{2}\|y-x\|^{2}.$
Taking $y=x^{*}$, the minimizer of $f$, we have
$f(x^{*})\ \leq\ f(x)+\langle\nabla f(x),\,x^{*}-x\rangle+\frac{L}{2}\|x-x^{*}\|^{2},$
or equivalently,
$f(x)-f(x^{*})\ \geq\ \langle\nabla f(x),\,x-x^{*}\rangle-\frac{L}{2}\|x-x^{*}\|^{2}.$
By the Baillon–Haddad theorem —Bauschke & Combettes (2017), Corollary 18.16—
$\langle\nabla f(x),\,x-x^{*}\rangle\ \geq\ \frac{1}{L}\|\nabla f(x)\|^{2}.$
Moreover, by the gradient $L$-Lipschitz property,
$\|\nabla f(x)\|\ \leq\ L\|x-x^{*}\|.$
Substituting into the previous inequality, we deduce
$f(x)-f(x^{*})\ \geq\ \frac{1}{2L}\|\nabla f(x)\|^{2}.$
(b) Upper bound. If $f$ is $m$-strongly convex, then for all $x,y$,
$f(y)\ \geq\ f(x)+\langle\nabla f(x),\,y-x\rangle+\frac{m}{2}\|y-x\|^{2}.$
Taking $y=x^{*}$, we get
$f(x^{*})\ \geq\ f(x)+\langle\nabla f(x),\,x^{*}-x\rangle+\frac{m}{2}\|x-x^{*}\|^{2},$
which is equivalent to
$f(x)-f(x^{*})\ \leq\ \langle\nabla f(x),\,x-x^{*}\rangle-\frac{m}{2}\|x-x^{*}\|^{2}.$
Letting $g:=\nabla f(x)$ and $d:=x-x^{*}$, we maximize the quadratic form
$\psi(d):=\langle g,\,d\rangle-\frac{m}{2}\|d\|^{2}$
with respect to $d$. The maximizer is $d=g/m$, and the maximum value is
$\max_{d}\psi(d)=\frac{1}{2m}\|g\|^{2}.$
Therefore,
$f(x)-f(x^{*})\ \leq\ \frac{1}{2m}\|\nabla f(x)\|^{2}.$
As the dynamics of AdaGrad and RMSProp are highly similar, differing primarily in the structure of the kernel, we may state the convergence and stability theorem in a unified form for both cases, with the sole difference arising from the Borel measure —Definition 1. Hence,
Theorem 1.
Under Assumptions 3 and 4, the solutions of
$\dot{\theta}^{i}(t)\;=\;-\,\frac{g^{i}(t)}{\sqrt{M_{\nu}[g^{2}](t+\alpha)+\epsilon}}$ (21)
satisfy the bound
$\|\theta(t)-\theta^{*}\|\ \leq\ \|\theta(0)-\theta^{*}\|\,e^{-m\,\mu_{0}\,t},\qquad\forall t\geq 0,$ (22)
where $\theta^{*}$ is the unique minimizer of $f$, $g_{i}(t):=\nabla_{i}f(\theta(t))$, $M_{\nu}[\cdot]$ is the memory operator with kernel $K_{\nu}$ defined in Definition 1, and